An epicyclic train is shown in Fig. 13.42. Internal gear A is keyed to the driving shaft and has 30 teeth. Compound wheel C and D of 20 and 22 teeth respectively are free to rotate on the pin fixed to the arm P which is rigidly connected to the driven shaft. Internal gear B which has 32 teeth is fixed. If the driving shaft runs at 60 r.p.m. clockwise, determine the speed of the driven shaft. What is the direction of rotation of driven shaft with reference to driving shaft?



1
Expert's answer
2021-06-04T14:01:03-0400

As the gear is fixed, "\implies \frac{33x}{32}+y=0 \implies \frac{33x}{32}=-y"

x+y=60

So, "\frac{33x}{32}-x=-60 \implies \frac{x}{32}=-60 \implies x=-1920 rpm"

Speed of the driven shaft = Arm speed

"y=-\frac{33x}{32}=\frac{33*1920}{32}=1980 rpm" clockwise direction